OpenStudy (anonymous):

Find two positive numbers whose difference is 8 and the sum of whose squares is 914. Give the answers in ascending order

OpenStudy (anonymous):

i just need the equations

OpenStudy (anonymous):

would it be x-y=8 for the first one?

OpenStudy (anonymous):

Yes, and x^2+y^2=914 for the second.

OpenStudy (anonymous):

thank you, i tried that before, but i didn't get the right answer

OpenStudy (anonymous):

can i add y^2 + y^2, and it'll be y^4?

OpenStudy (anonymous):

i'm getting y^2 +y^2=914-64

OpenStudy (anonymous):

but then i'm not sure what to do after; do i add the y's or do i take the square of 850?

OpenStudy (anonymous):

If you do y^2+y^2, it will be 2y^2 Then 2y^2=850, then divide by 2--- y^2=425 But that is not a perfect square. Hole on. Let me try something.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x-y=8 ----> -y=8-x---->y=-8+x k=x^2+y^2 k=x^2+(-8+x)^2 k=x^2+64-16x+x^2 k=2x^2-16x+64 Next we factor. Before I continue, did you understand that part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it'll be 2(x -8)(x-8)

OpenStudy (anonymous):

and then x-8=0 x=8 and same thing for the other x?

OpenStudy (anonymous):

Yes, that will be right, but now, I'm trying to figure how we will find what y is.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yay I got the answer. So I did something wrong. Instead of k=2x^2-16x+64 it should have been 914=2x^2-16x+64 Then subtract 914 from both sides. 0=2x^2-16x-850 Then divide the right side by 2 0=x^2-8x-425 Factor 0=(x-25)(x+17) Then solve x=25 and x=17 25-17=9 625+289=914 So the numbers are 25 and 17. That took awhile lol

OpenStudy (anonymous):

LOL yes that's right! Thank you so much!

OpenStudy (anonymous):

Np. It would've bugged me all day if I didn't figure it out. :)

OpenStudy (anonymous):

i'm sure it would have! lol

OpenStudy (anonymous):

are you also good at logs?

OpenStudy (anonymous):

Yes. Do you have a log question?

OpenStudy (anonymous):

yes i do, i'll post it now :)